OpenStudy (anonymous):

If f(x) = 7x2 - 8, what is f(2)?

OpenStudy (anonymous):

the question it self

OpenStudy (anonymous):

f(2) does equal f(x) = 7x2 - 8

OpenStudy (driftracer305):

lol what is 28-8

OpenStudy (vishweshshrimali5):

If you are given f(x) and you want to find f(c) where c is some number/constant, then all you have to do is to put c in place of x in the original expression of f(x) and simplify it. For example, \[\large{f(x) = 7x^2 - 8}\] If I want to find f(1) Then, put 1 in place of x in original expression of f(x): \[\large{f(1) = 7*(1)^2 - 8}\] Simplify it: \[\large{f(1) = 7 - 8 = -1}\] This is would be the value Similarly you can find out f(2)